# SEMF: Supervised Expectation-Maximization Framework for Predicting Intervals

### v3

@@ Appendix B Datasets for tabular benchmark
Table 4: Summary of benchmark tabular datasets retained from [45]
@@
Appendix C Optimal set of hyper-parameters
The hyper-parameter tuning for SEMF is implemented and monitored using Weights & Biases [66]. A random search is done in the hyper-parameter space for a maximum of 500 iterations on all 11 datasets, focusing on tuning the models only on the complete datasets. Key hyper-parameters are varied across a predefined set to balance accuracy and computational efficiency. The following grid is used for hyper-parameter tuning: the number of importance sampling operations $R\in\{5,10,25,50,100\}$ (100 is omitted for MultiMLPS), nodes per latent dimension $m_{k}\in\{1,5,10,20,30\}$, and standard deviations $\sigma_{m_{k}}\in\{0.001,0.01,0.1,1.0\}$. Early stopping steps (patience) are set to five or ten, and $R_{\text{infer}}$ is explored at [30, 50, 70]. The option to run the models in parallel must be consistently enabled. Table 5 shows the optimal set of hyper-parameters. This table includes common hyper-parameters for complete and 50% datasets and another part showing $\xi_{\mathrm{nodes}}$, which is tuned manually and only relevant to the missing data.
Table 5: Hyper-parameters for MultiXGBs, MultiETs, and MultiMLPs used for both complete and missing data.
Dataset Complete and Missing Missing
R $m_{k}$ $\sigma_{k}$ Patience $R_{\mathrm{infer}}$ $\xi_{\mathrm{nodes}}$
MultiXGBs
space_ga 10 30 1.0 5 70 100
cpu_activity 5 30 1.0 5 70 50
naval_propulsion_plant 5 30 0.01 5 50 100
miami_housing 5 10 0.1 5 50 50
kin8nm 5 30 1.0 10 70 100
concrete_compressive_strength 25 30 1.0 5 70 100
cars 50 10 1.0 10 70 100
energy_efficiency 5 1 0.01 10 70 100
california_housing 5 10 0.1 10 50 100
airfoil_self_noise 25 1 0.01 10 70 100
QSAR_fish_toxicity 50 30 1.0 5 70 100
MultiETs
space_ga 10 30 1.0 10 70 100
cpu_activity 5 30 1.0 10 70 50
naval_propulsion_plant 5 30 0.01 10 50 100
miami_housing 10 10 0.1 10 50 50
kin8nm 5 30 1.0 10 70 100
concrete_compressive_strength 25 30 1.0 10 70 100
cars 100 5 0.1 10 100 100
energy_efficiency 5 1 0.01 10 70 100
california_housing 5 10 0.1 5 50 100
airfoil_self_noise 25 1 0.01 10 70 100
QSAR_fish_toxicity 50 30 1.0 10 70 100
MultiMLPs
space_ga 25 10 0.001 10 50 100
cpu_activity 5 20 0.001 5 50 50
naval_propulsion_plant 5 20 0.001 5 50 100
miami_housing 5 20 0.01 5 50 50
kin8nm 5 20 0.001 5 50 100
concrete_compressive_strength 5 30 0.001 10 50 100
cars 5 30 0.1 5 50 100
energy_efficiency 50 30 0.1 10 50 100
california_housing 5 20 0.01 5 50 100
airfoil_self_noise 25 10 0.01 10 50 100
QSAR_fish_toxicity 50 30 1.0 10 70 100
MultiXGBs and MultiMLPs benefit from early stopping to reduce computation time in complete and incomplete cases. Similarly, the baseline models for these instances use the same hyper-parameters for early stopping. Further, the number of epochs in the case of MultiMLPs is set as 1000, except for energy_efficiency and QSAR_fish_toxicity, where this is changed to 5000. Any model-specific hyperparameter we did not specify in this paper remains at the implementation’s default value (e.g., the number of leaves in XGBoost from [50]). Along with the supplementary code, we provide three additional CSV files: one for the results and hyperparameters of all 330 runs and the other two for the optimal hyperparameters of SEMF models, both raw (directly from SEMF) and conformalized.
Additional conditions are applied only for experiments with missing data. Datasets california_housing, cpu_activity, miami_housing, and naval_propulsion_plant—have a patience of five to expedite the training process. Additionally, for california_housing and cpu_activity, the $R_{\text{infer}}$ value is set to 30, while for all the other datasets, it is set to 50. We do this to ensure efficient computation, speed, and memory usage (especially for the GPU).
For training MultiXGBs and MultiETs, the computations are performed in parallel using CPU cores (Intel® Core i9-13900KF). For MultiMLPs, they are done on a GPU (NVIDIA® GeForce RTX 4090). The GPU is also consistently used for the missing data simulator and training $p_{\xi}$. All the computations are done on a machine with 32 GB of memory. The code provides further details on hardware and reproducibility.
Appendix D Metrics for prediction intervals
D.1 Common metrics
The most common metrics for evaluating prediction intervals [52, 53] are:
@@ D.1 Common metrics
D.2 Impact of relative metrics for modeling
@@
Appendix E Full conformalized results
@@
Interval Predictions Point Predictions
Relative Absolute Relative Absolute
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
space_ga 2% -2% 4% 0.95 2.40 0.26 -10% -13% 0.60
space_ga 10% -1% 10% 0.95 2.34 0.25 -10% -10% 0.62
space_ga 8% -2% 9% 0.93 2.24 0.24 -9% -8% 0.60
space_ga 6% -2% 7% 0.94 2.38 0.26 -8% -10% 0.59
space_ga 7% -1% 7% 0.95 2.55 0.28 -8% -10% 0.59
cpu_activity 7% 1% 5% 0.95 0.49 0.09 22% 2% 0.98
cpu_activity 20% 1% 17% 0.93 0.45 0.09 20% -5% 0.98
cpu_activity 20% 0% 17% 0.93 0.45 0.09 21% 0% 0.98
cpu_activity 21% 1% 16% 0.94 0.45 0.09 23% -2% 0.98
cpu_activity 10% 2% 7% 0.95 0.47 0.09 20% -3% 0.98
naval_propulsion_plant 163% -1% 62% 0.95 0.36 0.11 -15% -12% 0.99
naval_propulsion_plant 190% 0% 66% 0.95 0.37 0.11 -20% -14% 0.99
naval_propulsion_plant 151% 0% 60% 0.96 0.40 0.12 -12% -9% 0.99
naval_propulsion_plant 185% -1% 65% 0.94 0.35 0.10 -9% -15% 0.99
naval_propulsion_plant 169% -1% 63% 0.96 0.37 0.11 -15% -11% 0.99
miami_housing -2% 0% -2% 0.95 0.99 0.12 4% 1% 0.91
miami_housing -10% 0% -10% 0.95 1.03 0.13 7% 7% 0.90
miami_housing -6% 1% -8% 0.95 0.99 0.12 9% 5% 0.91
miami_housing -7% 0% -8% 0.95 0.99 0.12 3% -1% 0.91
miami_housing -10% 0% -12% 0.95 1.05 0.13 -4% 0% 0.90
kin8nm 6% 0% 6% 0.94 2.27 0.45 -21% -24% 0.62
kin8nm 4% 1% 3% 0.95 2.39 0.47 -19% -21% 0.64
kin8nm 9% 1% 8% 0.94 2.20 0.43 -17% -20% 0.64
kin8nm 9% 1% 7% 0.95 2.36 0.46 -21% -22% 0.63
kin8nm 3% 1% 2% 0.94 2.33 0.46 -20% -23% 0.62
concrete_compressive_strength 40% -3% 31% 0.93 1.42 0.29 -16% -22% 0.86
concrete_compressive_strength 47% -6% 36% 0.91 1.48 0.30 -29% -52% 0.84
concrete_compressive_strength 59% -2% 39% 0.96 1.60 0.32 -21% -34% 0.86
concrete_compressive_strength 35% -3% 28% 0.95 1.48 0.30 -28% -39% 0.83
concrete_compressive_strength 24% 0% 19% 0.97 1.59 0.32 -34% -52% 0.86
cars 56% -4% 38% 0.91 0.70 0.12 -7% -4% 0.95
cars 36% -5% 30% 0.90 0.86 0.14 -4% -4% 0.95
cars 30% -3% 25% 0.91 0.76 0.13 2% 5% 0.96
cars 65% -1% 40% 0.93 0.75 0.13 -5% -1% 0.95
cars 16% -4% 17% 0.91 0.87 0.15 -2% 0% 0.95
energy_efficiency 165% -5% 64% 0.92 0.20 0.06 -50% -29% 1.00
energy_efficiency 288% 0% 74% 0.91 0.14 0.04 -3% -13% 1.00
energy_efficiency 217% -1% 69% 0.96 0.18 0.05 -21% -31% 1.00
energy_efficiency 253% -3% 73% 0.92 0.16 0.04 17% 16% 1.00
energy_efficiency 185% -10% 68% 0.91 0.23 0.07 -17% -22% 1.00
california_housing -5% 0% -6% 0.95 1.82 0.43 -2% -3% 0.81
california_housing 4% 0% 4% 0.95 1.72 0.41 -1% -2% 0.81
california_housing -4% 1% -5% 0.95 1.76 0.42 0% -1% 0.82
california_housing 1% 0% 1% 0.95 1.74 0.41 0% 0% 0.81
california_housing -3% 1% -4% 0.95 1.79 0.42 -1% -1% 0.82
airfoil_self_noise 15% -4% 16% 0.95 1.73 0.37 -82% -93% 0.86
airfoil_self_noise 45% 1% 30% 0.97 1.42 0.30 -18% -18% 0.93
airfoil_self_noise 5% 1% 4% 0.98 1.86 0.40 -73% -76% 0.85
airfoil_self_noise -6% 0% -7% 0.98 2.15 0.46 -109% -138% 0.78
airfoil_self_noise 46% -2% 33% 0.95 1.37 0.29 -41% -43% 0.89
QSAR_fish_toxicity 21% -2% 19% 0.88 2.12 0.33 3% 3% 0.53
QSAR_fish_toxicity 39% -9% 35% 0.85 1.97 0.31 8% 7% 0.58
QSAR_fish_toxicity 43% -3% 32% 0.89 2.21 0.34 4% 3% 0.57
QSAR_fish_toxicity 38% -3% 30% 0.89 2.11 0.33 -2% -1% 0.54
QSAR_fish_toxicity 29% -6% 28% 0.85 2.11 0.33 0% -6% 0.52

### v4

@@ Appendix B Datasets for tabular benchmark
Table 6: Summary of benchmark tabular datasets retained from [44]
Appendix C Metrics for prediction intervals
C.1 Metrics’ definitions
The most common metrics for evaluating prediction intervals [58, 59, 60, 61] are:
@@ C.1 Metrics’ definitions
  • Continuous Ranked Probability Score (CRPS): We can approximate a full predictive cumulative distribution function by assuming that the forecast is uniformly distributed between its lower and upper bounds. Let $L$ and $U$ denote the lower and upper bounds (corresponding to the $\alpha/2$ and $21-\alpha/$ 2 quantiles, respectively). Then, for an observation $y$, we define
    $\text{CRPS}(y,L,U)=\begin{cases}L-y+\frac{U-L}{3},&y\leq L,\\ \frac{(y-L)^{3}+(U-y)^{3}}{3\,(U-L)^{2}},&y\in[L,U],\\ y-U+\frac{U-L}{3},&y\geq U.\end{cases}$ (23)
    This closed-form expression yields a lower CRPS when the forecast is both sharper and better calibrated.
  • Quantile (pinball) Loss: For a given quantile level $\tau\in(0,1)$ and prediction $q$, the pinball loss is defined as:
    $\ell_{\tau}(y,q)=(y-q)\left(\tau-\mathds{1}\{y\leq q\}\right),$ (24)
    where $\mathds{1}\{y\leq q\}$ equals 1 if $y\leq q$ and 0 otherwise. This loss penalizes underestimation and overestimation asymmetrically and is used to tune the quantile predictions.
C.2 Impact of relative metrics for modeling
@@
Appendix D Full synthetic experiments results
Relative Metrics Absolute Metrics
Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
Normal
0 17% 15% 6% 11% 0% 1.78 0.54 0.97
10 11% 11% 3% 5% -1% 1.60 0.60 0.96
20 19% 18% 8% 14% -2% 1.89 0.50 0.95
30 10% 10% 5% 10% -1% 1.99 0.48 0.96
40 35% 28% 10% 17% -3% 2.49 0.37 0.93
Uniform
0 -2% -5% -1% 0% 2% 2.36 0.42 0.98
10 34% 26% 12% 26% -1% 2.12 0.47 0.99
20 9% 8% 5% 12% 1% 2.11 0.47 0.99
30 28% 26% 9% 19% -5% 3.59 0.25 0.91
40 58% 39% 23% 32% -3% 3.61 0.26 0.95
Lognormal
0 23% 19% 15% 9% 0% 2.08 0.46 0.96
10 59% 39% 18% 7% -3% 2.07 0.46 0.95
20 8% 9% 8% 2% -1% 1.96 0.48 0.95
30 9% 8% 9% 5% 0% 1.59 0.62 0.98
40 38% 28% 15% 14% 0% 2.30 0.42 0.97
Gumbel
0 4% 1% 7% 16% 4% 1.61 0.60 0.98
10 17% 18% 7% 4% -3% 1.78 0.53 0.94
20 20% 17% 14% 16% -1% 2.01 0.48 0.96
30 -5% -9% 1% 0% 4% 1.75 0.55 0.96
40 15% 14% 10% 5% -1% 2.02 0.47 0.95
Table 7: Test results for MultiXGBs, over 5 seeds, with cosine $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
Relative Metrics Absolute Metrics
Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
Normal
0 8% 8% 4% 0% -1% 1.98 0.48 0.96
10 17% 18% 6% 12% -3% 1.66 0.57 0.94
20 1% 0% 0% 2% 1% 1.81 0.54 0.97
30 15% 14% 8% 16% -1% 1.99 0.49 0.97
40 32% 27% 8% 9% -4% 2.46 0.38 0.94
Uniform
0 35% 28% 10% 21% -3% 2.80 0.35 0.97
10 3% -1% 3% 8% 4% 2.45 0.41 0.99
20 7% 5% 4% 11% 1% 2.58 0.37 0.96
30 9% 8% 6% 14% 0% 3.42 0.28 0.97
40 29% 25% 7% 12% -3% 3.63 0.26 0.96
Lognormal
0 -14% -18% 8% -10% 1% 1.85 0.52 0.96
10 18% 17% 10% -2% -2% 1.87 0.51 0.95
20 12% 12% 13% 0% -1% 2.03 0.47 0.96
30 -6% -8% 5% -4% 2% 1.74 0.56 0.98
40 14% 13% 11% 2% -2% 2.10 0.46 0.97
Gumbel
0 -4% -7% 9% 0% 2% 1.66 0.59 0.97
10 8% 7% 8% -2% 0% 1.72 0.56 0.97
20 -2% -2% 7% 5% 0% 1.90 0.50 0.96
30 2% 4% 6% -2% -2% 1.85 0.51 0.95
40 14% 14% 11% 5% -2% 1.96 0.50 0.97
Table 8: Test results for MultiETs, over 5 seeds, with cosine $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
Relative Metrics Absolute Metrics
Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
Normal
0 0% 1% -2% 0% -1% 1.98 0.48 0.95
10 9% 8% 4% 5% 0% 1.71 0.56 0.96
20 -2% -1% -1% -3% -1% 2.04 0.46 0.94
30 -1% -2% 1% 3% 0% 2.20 0.44 0.97
40 7% 7% 1% 5% 0% 2.48 0.39 0.96
Uniform
0 2% 2% 0% 8% 0% 2.98 0.33 0.98
10 6% 2% 5% 14% 3% 2.80 0.35 0.98
20 3% 7% 0% -5% -4% 3.02 0.31 0.94
30 -2% -2% -1% -4% 0% 3.85 0.25 0.96
40 13% 10% 8% 23% 1% 3.67 0.27 0.99
Lognormal
0 2% 3% 1% 4% 0% 1.71 0.57 0.98
10 18% 16% 7% 8% 0% 1.99 0.48 0.95
20 -4% -5% -2% 0% 1% 1.86 0.52 0.98
30 15% 13% 7% 9% -1% 1.74 0.56 0.98
40 4% 5% 1% 2% -1% 2.10 0.46 0.97
Gumbel
0 -3% -1% -2% -4% -1% 1.54 0.64 0.98
10 8% 8% 1% 4% -1% 1.88 0.51 0.95
20 8% 8% 3% 4% 0% 2.14 0.44 0.94
30 13% 14% 2% -4% -2% 1.87 0.51 0.95
40 -3% -5% -2% -2% 2% 1.76 0.56 0.98
Table 9: Test results for MultiMLPs, over 5 seeds, with cosine $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
Relative Metrics Absolute Metrics
Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
Normal
0 35% 26% 28% 16% 0% 5.80 0.17 0.96
10 65% 41% 34% 35% -3% 4.80 0.20 0.96
20 40% 29% 31% 16% 0% 4.41 0.21 0.94
30 30% 23% 26% 23% 0% 5.32 0.18 0.95
40 28% 22% 27% 28% 0% 5.83 0.16 0.95
Uniform
0 80% 44% 42% 30% 0% 10.12 0.09 0.94
10 54% 34% 47% 32% 1% 5.78 0.17 0.98
20 62% 38% 40% 31% 0% 5.62 0.17 0.96
30 62% 39% 39% 16% -1% 7.28 0.13 0.95
40 68% 41% 39% 18% -1% 7.93 0.12 0.97
Lognormal
0 2% 1% 20% 0% 1% 4.37 0.22 0.96
10 -2% -2% 14% -4% 1% 2.92 0.34 0.99
20 7% 2% 27% 6% 5% 3.81 0.25 0.94
30 -13% -16% 15% 3% 1% 3.28 0.29 0.96
40 7% 6% 23% 16% 0% 3.42 0.29 0.98
Gumbel
0 34% 27% 30% 20% -2% 5.03 0.19 0.96
10 49% 35% 24% 21% -3% 3.70 0.26 0.95
20 0% 0% 9% 3% 1% 2.81 0.35 0.99
30 11% 7% 17% 14% 4% 4.49 0.21 0.95
40 18% 15% 21% 20% 1% 4.37 0.22 0.97
Table 10: Test results for MultiXGBs, over 5 seeds, with quadratic $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
Relative Metrics Absolute Metrics
Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
Normal
0 7% 7% 0% -4% -1% 5.77 0.16 0.92
10 35% 26% 21% 23% 0% 4.66 0.21 0.98
20 29% 24% 15% 14% -2% 4.69 0.20 0.93
30 26% 21% 15% 13% -1% 4.91 0.19 0.95
40 21% 16% 18% 12% 1% 4.79 0.20 0.98
Uniform
0 28% 23% 16% -4% -1% 8.81 0.11 0.94
10 25% 20% 24% 21% 1% 4.95 0.20 0.99
20 47% 34% 25% 7% -3% 4.51 0.21 0.96
30 30% 22% 22% 15% 2% 5.85 0.17 0.98
40 71% 42% 33% 21% 0% 7.34 0.13 0.97
Lognormal
0 -9% -13% 0% -6% 3% 4.34 0.22 0.96
10 -1% -2% 13% 12% 1% 3.36 0.29 0.98
20 33% 27% 16% 10% -2% 4.09 0.23 0.95
30 8% 8% 11% 9% 0% 3.29 0.30 0.97
40 10% 8% 15% 9% 1% 3.78 0.26 0.97
Gumbel
0 5% 4% 10% 7% 1% 4.67 0.21 0.98
10 42% 30% 27% 28% -1% 4.55 0.21 0.94
20 33% 25% 16% 15% -1% 3.44 0.28 0.97
30 -3% -4% 5% 0% 1% 3.86 0.25 0.97
40 68% 42% 30% 24% -2% 3.97 0.24 0.97
Table 11: Test results for MultiETs, over 5 seeds, with quadratic $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
Relative Metrics Absolute Metrics
Seed $\Delta$CWR $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball $\Delta$PICP CWR NMPIW PICP
Normal
0 -14% -20% 3% 6% 3% 6.63 0.15 0.97
10 24% 21% 13% 19% -2% 6.07 0.16 0.96
20 0% -4% 2% 6% 3% 5.81 0.16 0.93
30 -1% -5% 7% 17% 4% 6.13 0.16 0.95
40 15% 14% 9% 20% 0% 7.47 0.13 0.95
Uniform
0 42% 26% 25% 36% 6% 15.86 0.06 0.97
10 27% 20% 14% 30% 2% 9.42 0.11 1.00
20 66% 41% 27% 29% -2% 10.02 0.10 0.95
30 34% 26% 13% 9% 0% 12.50 0.08 0.96
40 15% 13% 16% 40% -1% 13.66 0.07 0.97
Lognormal
0 1% 2% 4% 0% -1% 5.41 0.18 0.97
10 3% 2% 8% 0% 1% 5.23 0.18 0.96
20 10% 11% 3% -10% -2% 4.79 0.20 0.95
30 -12% -15% 4% 0% 1% 3.86 0.25 0.97
40 4% 3% 2% 4% 0% 4.88 0.20 0.97
Gumbel
0 38% 30% 19% 18% -3% 6.90 0.14 0.96
10 32% 25% 23% 23% -1% 4.56 0.21 0.96
20 -6% -8% -3% -5% 2% 4.10 0.23 0.96
30 13% 12% 3% 8% -1% 5.40 0.18 0.96
40 -3% -5% 10% 10% 1% 5.14 0.19 0.96
Table 12: Test results for MultiMLPs, over 5 seeds, with quadratic $f(x)$ using 2 predictors, organized by distribution and seed. Relative metrics are in bold when positive.
Appendix E Full conformalized results on benchmark datasets
Relative Metrics Absolute Metrics
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball CWR NMPIW PICP
space_ga 9% 0% 9% 12% 10% 6.28 0.15 0.95
space_ga 5% 2% 2% 11% 13% 2.95 0.32 0.95
space_ga 4% 1% 2% 10% 9% 3.23 0.30 0.97
space_ga 21% -2% 19% 11% 9% 3.37 0.28 0.96
space_ga 16% -2% 15% 11% 6% 6.74 0.14 0.94
cpu_activity 7% 0% 7% 17% 10% 9.84 0.10 0.96
cpu_activity -3% 1% -4% 10% 0% 8.69 0.11 0.96
cpu_activity 11% -1% 11% 14% 10% 10.20 0.09 0.94
cpu_activity 11% 1% 10% 26% 36% 10.14 0.09 0.95
cpu_activity 1% 1% 0% 13% 18% 9.99 0.10 0.95
naval_propulsion_plant 200% -3% 68% 73% 43% 9.19 0.10 0.92
naval_propulsion_plant 114% 0% 53% 66% 31% 7.21 0.13 0.95
naval_propulsion_plant 159% 1% 61% 74% 50% 7.97 0.12 0.96
naval_propulsion_plant 187% 0% 65% 74% 50% 8.73 0.11 0.95
naval_propulsion_plant 156% 2% 60% 71% 50% 8.24 0.12 0.96
miami_housing 9% -1% 9% 14% -5% 9.08 0.10 0.95
miami_housing -4% 0% -4% 8% -12% 8.28 0.12 0.96
miami_housing 2% 0% 3% 9% -19% 8.49 0.11 0.94
miami_housing 12% 0% 10% 17% -5% 9.23 0.10 0.95
miami_housing 8% -1% 9% 11% -21% 8.70 0.11 0.95
kin8nm 16% -1% 15% 22% 3% 2.29 0.41 0.94
kin8nm 19% -1% 16% 25% 11% 2.31 0.41 0.94
kin8nm 16% -1% 14% 20% 6% 2.27 0.42 0.95
kin8nm 19% -1% 17% 22% 8% 2.29 0.41 0.95
kin8nm 19% -1% 17% 24% 11% 2.40 0.40 0.95
concrete_compressive_strength 31% -1% 24% 35% 25% 2.95 0.33 0.96
concrete_compressive_strength 73% -5% 45% 34% 14% 3.60 0.26 0.92
concrete_compressive_strength 27% -3% 24% 29% 6% 3.14 0.29 0.92
concrete_compressive_strength 47% -2% 34% 33% 9% 3.14 0.30 0.94
concrete_compressive_strength 24% -1% 20% 27% 4% 3.25 0.29 0.93
cars 32% -2% 25% 25% 6% 5.78 0.17 0.95
cars 10% -1% 10% 20% 18% 5.02 0.19 0.95
cars 27% 0% 22% 26% 17% 5.08 0.19 0.95
cars 67% -3% 42% 42% 32% 6.15 0.15 0.94
cars 22% -2% 20% 23% 15% 4.63 0.21 0.97
energy_efficiency 114% 1% 52% 66% 43% 15.29 0.06 0.96
energy_efficiency 112% 0% 53% 67% 50% 14.18 0.07 0.95
energy_efficiency 109% -2% 53% 64% 44% 10.78 0.09 0.96
energy_efficiency 230% 8% 68% 81% 73% 17.66 0.05 0.96
energy_efficiency 72% 5% 39% 59% 58% 11.84 0.08 0.95
california_housing 3% 0% 3% 23% -3% 2.09 0.46 0.96
california_housing 4% -1% 4% 18% -10% 2.27 0.42 0.94
california_housing 2% 0% 1% 21% -7% 2.29 0.41 0.95
california_housing 0% 0% 0% 15% -9% 2.27 0.41 0.94
california_housing -1% 0% -1% 15% -10% 2.28 0.42 0.95
airfoil_self_noise -5% 0% -5% 9% -19% 2.60 0.37 0.96
airfoil_self_noise 6% 0% 6% 20% 4% 2.56 0.38 0.98
airfoil_self_noise -32% 0% -47% -36% -43% 1.57 0.63 0.99
airfoil_self_noise -4% 0% -5% 8% -10% 2.26 0.43 0.97
airfoil_self_noise -8% -1% -7% 9% -15% 2.43 0.39 0.95
QSAR_fish_toxicity 2% 0% 1% 8% 10% 1.80 0.54 0.97
QSAR_fish_toxicity 11% -1% 10% 8% 12% 1.63 0.59 0.96
QSAR_fish_toxicity 4% 2% 3% 9% 8% 1.80 0.54 0.97
QSAR_fish_toxicity 26% 0% 20% 17% 18% 1.55 0.64 0.99
QSAR_fish_toxicity 53% -1% 35% 24% 30% 1.73 0.57 0.98
@@ Appendix E Full conformalized results on benchmark datasets
Relative Metrics Absolute Metrics
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball CWR NMPIW PICP
space_ga 4% -2% 6% 6% -9% 5.86 0.16 0.95
space_ga -2% 0% -2% 7% -3% 2.66 0.36 0.96
space_ga 10% -1% 10% 7% -5% 3.19 0.30 0.97
space_ga -2% 0% -2% 7% -5% 2.99 0.33 0.97
space_ga 15% -3% 16% 11% -9% 7.22 0.13 0.91
cpu_activity 5% -5% 10% 2% -22% 8.69 0.11 0.94
cpu_activity 1% -3% 4% 7% -11% 8.05 0.12 0.96
cpu_activity 3% -5% 8% 5% -11% 8.26 0.11 0.94
cpu_activity -1% -4% 2% -2% -22% 8.14 0.12 0.95
cpu_activity 3% -4% 7% 2% -22% 8.11 0.12 0.95
naval_propulsion_plant 154% -6% 63% 69% 40% 3.55 0.27 0.94
naval_propulsion_plant 153% 0% 60% 68% 46% 3.95 0.24 0.95
naval_propulsion_plant 160% 0% 61% 72% 52% 3.86 0.25 0.96
naval_propulsion_plant 136% 0% 58% 66% 41% 3.67 0.26 0.95
naval_propulsion_plant 118% -4% 56% 64% 41% 3.23 0.29 0.95
miami_housing -13% 1% -16% -4% -75% 6.25 0.15 0.96
miami_housing -6% 0% -6% 0% -67% 6.71 0.14 0.95
miami_housing -5% 1% -6% 0% -67% 6.47 0.15 0.95
miami_housing -5% 0% -4% -1% -75% 6.75 0.14 0.95
miami_housing -15% 1% -19% -7% -81% 6.03 0.16 0.95
kin8nm 8% 0% 7% 15% 0% 2.14 0.44 0.95
kin8nm 7% 0% 7% 13% -3% 2.11 0.45 0.95
kin8nm 7% 0% 7% 14% 0% 2.10 0.46 0.96
kin8nm 5% 1% 4% 12% 0% 2.08 0.46 0.95
kin8nm 10% 1% 8% 17% 0% 2.22 0.43 0.95
concrete_compressive_strength 26% -3% 23% 12% 0% 3.27 0.29 0.94
concrete_compressive_strength 5% 0% 5% 8% 0% 2.97 0.33 0.97
concrete_compressive_strength 13% 3% 9% 12% 3% 3.39 0.28 0.93
concrete_compressive_strength -14% 2% -20% 4% -12% 2.73 0.35 0.95
concrete_compressive_strength 3% 0% 3% 9% -4% 2.87 0.34 0.97
cars 12% -1% 12% 3% 0% 4.95 0.19 0.96
cars 2% -2% 4% -1% 0% 4.97 0.19 0.96
cars 29% -3% 24% 15% 15% 3.64 0.27 0.97
cars -25% 4% -40% -8% -5% 6.19 0.15 0.92
cars 54% -2% 36% 20% 28% 4.70 0.21 0.97
energy_efficiency 26% -3% 23% 12% 0% 15.20 0.06 0.96
energy_efficiency 16% -3% 17% 6% 0% 16.14 0.06 0.94
energy_efficiency 24% 0% 20% 9% 25% 14.43 0.07 0.99
energy_efficiency 8% 0% 8% 4% 0% 20.70 0.04 0.91
energy_efficiency 1% -1% 1% -15% -80% 13.19 0.07 0.96
california_housing -$-1$1% 16% -26% 1.55 0.62 0.96
california_housing -1% -2% 1% 20% -24% 1.68 0.57 0.95
california_housing 2% -2% 4% 24% -17% 1.70 0.56 0.95
california_housing 4% -3% 7% 20% -25% 1.76 0.53 0.93
california_housing 9% -3% 11% 25% -17% 1.83 0.52 0.94
airfoil_self_noise -21% -1% -24% -26% -40% 2.15 0.45 0.97
airfoil_self_noise 22% 0% 18% 19% 18% 2.91 0.34 0.98
airfoil_self_noise -20% -2% -22% -21% -36% 2.17 0.44 0.96
airfoil_self_noise -39% 1% -66% -52% -61% 1.52 0.64 0.98
airfoil_self_noise -6% -1% -5% -3% -12% 2.49 0.39 0.97
QSAR_fish_toxicity -13% 0% -14% -4% -10% 1.69 0.57 0.97
QSAR_fish_toxicity -20% 0% -26% -16% -19% 1.36 0.72 0.98
QSAR_fish_toxicity 17% -4% 18% 6% -2% 2.00 0.47 0.94
QSAR_fish_toxicity 17% -2% 16% 4% 2% 1.87 0.52 0.97
QSAR_fish_toxicity -18% 0% -22% -7% -14% 1.60 0.61 0.97